Session: Feb 2011 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Mariotti at Thu Feb 24 15:21:15 2011) Success:
no known problems

Westerbork (Geert Kuper at Thu Feb 24 18:22:33 2011) Success:
No Known problems

Onsala (Michael Lindqvist at Thu Feb 24 20:32:24 2011) Success:
Apart from RFI, no known problems.

Urumqi (Wenjun Yang at Fri Feb 25 02:32:07 2011) Success:
No P-cal signal to be found because of the cable connect point is not good, it is ok in Exp. ey010d.

Urumqi (Wenjun Yang at Fri Feb 25 02:32:39 2011) Success:
It is heavy snow.

Hartebeesthoek (Jonathan Quick at Fri Feb 25 08:11:13 2011) Success: RFI
No problems apart from RFI in VCs 7 and 8

Torun (K. Borkowski at Fri Feb 25 13:05:13 2011) Minor failures:
The ftp-files would not be produced by Mark5A. A number of correlation tries of manually extracted disk2file data failed supposedly because we could not hit proper timestamp (Mark5A reported it cannot recognize the beginnings of scans). Basing on this experience, one improvement in SCHED could be suggested. Presently the .skd or .vex file shows seconds into a scan as e.g. 753 to 755 for 2-second ftp data. Now we were not clear whether the seconds are 754 and 755 or 753 and 754. Thus, it would be less ambiguous if the numbers indicated the first and the last second to be extracted (ftp-ed), so that second 1 would indicate the very beginning of a scan. Or one might have the seconds given as real numbers, e.g. 753.0 and 755.0, then the very beginning of a scan would be numbered as 0.0.

Effelsberg (Uwe Bach at Tue Mar 01 14:28:10 2011) Failure:
One of the MK4 cables that connect formater and Mark5A was connected to the VLBA port. It was the "Set 2, odd" connector. Don't know if that helps to correlate at least some of the data or if everything is lost. Should be solved for the user experiments. 27.2. Just noticed that the "Set 1, odd" attenuator on the ouput port wasn't connected properly. If this affects the recording, it might have some effect on the whole 18cm data until EM080C :-(

Shanghai (Bo Xias at Wed Mar 02 04:57:16 2011) Success:

Jodrell_Bank (Alastair Gunn at Tue Mar 29 11:24:44 2011) Minor failures: RECO
Briefly changed disk pack at 14:29 for testing. No other known problems.